Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for June 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, June 24, 1998

FOR DATES: From 7:00 a.m. Wednesday, June 17, 1998 to 7:00 a.m. Wednesday, June 24, 1998.

The Thermal Destruction Facility (TDF) was on-line for 149.58 hours and processed 4768.1 tons of contaminated soil during the period of this report. The TDF was on-line 89% of the time with an average hourly feed rate of 31.9 tons per hour while on-line. The contractor continues to work 24 hours a day, 7 days a week operating the TDF.

To date, the contractor has successfully treated approximately 60,489 tons of contaminated soil through the TDF amounting to 24% of the overall contaminated soil on-site.

During the time frame of this report, feed stopped several times due to a build up of soil materials on the dual feed screws. Feed was manually stopped several times as a precaution during local lightning storm activity. As the week progressed, feed rates were reduced, and the incidents of feed screw stoppages became more common. The contractor decided to replace the worn feed screws, the feed screw cooling jacket, and perform needed maintenance and cleaning of the TDF. On Tuesday, June 23, the TDF was placed in hot standby to perform routine maintenance and to perform TDF interlock inspections. The TDF is scheduled to be back on-line by Friday, June 26 or Saturday, June 27.

OHM's subcontractor, AirRecon, successfully completed stack testing for dioxins and furans, and metals and particulates on Thursday, June 18. USACE's subcontractor, Focus Environmental, observed stack sampling.

Construction activities during this period included wastewater management, excavation of soils within the exclusion zone, rock washing and decontamination of inorganic debris for backfilling, and handling of drums. The contractor continued the placement of Concover on the stockpile and placement of odor suppressant foam on the excavation face within the exclusion zone, and continued off-site and perimeter air monitoring.

During the time frame of this report, 2 drums with contents were excavated, sampled, tagged, and over packed. The over packed drums are currently in temporary storage within the exclusion zone. The contractor received finger print lab analyses (characterization) of previously overpacked drums. Excavation within the exclusion zone uncovered a large amount of contaminated demolition debris and drum carcasses. The demolition debris will be separated and the drum carcasses will be shipped to a hazardous waste landfill.

Related site activities during this period included:

* Weekly progress meeting held on Tuesday, June 16, 1998. * Junior partnering meeting of the Drake Cleanup Team held on Friday, June 19, 1998. * AirRecon conducted stack testing on June 16, 17, and18, 1998.

Major activities for the next week include: Restarting feed and operating the TDF, soil feed preparation, off-site and perimeter air monitoring, WWTP operations, ash handling, rock & inorganic debris handling, and drum handling and bulking.

PREPARED BY: Steve Gibson & David Modricker, USACE


Wednesday, June 17, 1998

FOR DATES: From 7:00 a.m. Wednesday, June 10, 1998 to 7:00 a.m. Wednesday, June 17, 1998.

The Thermal Destruction Facility (TDF) was on-line for approximately 153.45 hours and processed approximately 6413.2 tons of contaminated soil during the period of this report. The contractor continues to work 24 hours a day, 7 days a week operating the TDF.

To date, the contractor has successfully treated and backfilled 55,721 tons of contaminated soil amounting to 22% of the contaminated soil on-site.

During the time frame of this report, feed was stopped several times to remove a build up soil on the feed screws and as a precaution during lightning storm events. On Wednesday, June 10, an unplanned TRV opening occurred due to a local grid power failure. The TRV remained open for less than 1 minute. USACE personnel were informed by PPL that the power failure was probably caused by a biological interruption (squirrel) and could not have been anticipated.

Construction activities during the time frame of this report included feed preparation, debris separation, wastewater management, lime treatment of feed soils, rock washing, and decontamination of inorganic debris for backfilling. The contractor continued the placement of Concover on the stockpile, the placement of odor suppressant foam on the excavation face within the exclusion zone, and continued off-site and perimeter air monitoring activities. AIR-RECON arrived on-site and is currently performing stack testing for metals and particulates, and dioxins and furans. Stack testing is scheduled to finish on Thursday, June 18, 1998.

Related site activities during this period included:
* Weekly progress meeting held on Tuesday, June 16, 1998.
* Visit and tour of facility by Bob Carnein and Lock Haven University Students.

Major activities for the next week include: Operating the TDF, soil feed preparation, off-site and perimeter air monitoring, WWTP operations, ash handling, rock & inorganic debris washing, and drum bulking for shipment off-site.

PREPARED BY: Steve Gibson & David Modricker, USACE


Wednesday, June 10, 1998

FOR DATES: From 7:00 a.m. Wednesday, June 3, 1998 to 7:00 a.m. Wednesday, June 10, 1998.

The Thermal Destruction Facility (TDF) was on-line for approximately 135 hours and processed approximately 5339.8 tons of contaminated soil during the period of this report. The contractor continues to work 24 hours a day, 7 days a week operating the TDF.

To date, the contractor has successfully treated approximately 49,308 tons of contaminated soil through the TDF and all ash backfilled during the operations phase has met project requirements and specifications for ash placement.

During the time frame of this report, feed was stopped for an extended period of time to perform maintenance on the TDF. The TDF remained in hot standby while repairs were made to the lower dump valve that is located between the dry and the wet ash conveyors. In addition to the lower dump valve repairs, the feed screw cooling jacket was converted from a water cooled configuration to an air cooled configuration. Cooling configurations had to be switched because feed was becoming saturated with water from a crack in the cooling jacket while entering the rotary kiln. Once repairs had been made and the kiln reached operating temperature, feed was resumed on Tuesday at 1432 hours.

Construction activities during this period included wastewater management, rock washing and decontamination of inorganic debris for backfilling, feed preparation, debris separation, lime treatment of feed, ash handling, ash backfilling, placement of Concover on the stockpile within the exclusion zone, and continuation of off-site and perimeter air monitoring activities. During the time frame of this report, excavation activities within the exclusion zone were postponed due to equipment downtime, lack of additional stockpile area, accommodations to material handling in the feed preparation building, and to facilitate handling the large amount of site debris. Drum carcasses from the previous week's excavation and used PPE were shipped to a hazardous waste landfill.

Related site activities during this period included:

* Weekly progress meeting held on Tuesday, June 9, 1998. * A site tour was given to Lock Haven University students. * A field manager partnering meeting between the USACE, OHM, and PADEP was held to address staff concerns at the Drake Chemical Superfund Site. * Stack sampling for dioxin and furans, metals and particulates is scheduled for June 16th, 17th, and 18th.

Major activities for the next week include: Operating the TDF, excavation and drum handling, soil feed preparation, off-site and perimeter air monitoring, WWTP operations, ash handling, rock & inorganic debris washing, and drum handling. Stack sampling will be performed as stated above.

PREPARED BY: Steve Gibson & David Modricker, USACE


Wednesday, June 3, 1998

FOR DATES: From 7:00 a.m. Wednesday, May 27, 1998 to 7:00 a.m. Wednesday, June 3, 1998.

The Thermal Destruction Facility (TDF) was on-line for approximately 154.73 hours and processed approximately 6460.4 tons of contaminated soil during the time frame of this report. During this period, site soils were being processed at a rate of 41.7 tons per hour. The contractor continues to work 24 hours a day, 7 days a week operating the TDF.

To date, the contractor has successfully treated approximately 54,329 tons of contaminated soil through the TDF. All ash backfilled to date has met project requirements and specifications for backfill criteria.

As a precaution to power failure, feed was stopped several times during the time frame of this report because of local electrical storm events. Feed was resumed after lightning activity was observed to leave the area. A local power failure at 2045 hours on Sunday, May 31, caused an unscheduled Thermal Relief Valve (TRV) opening. Lightning had been observed in the distance and feed to the kiln had been stopped for 16 minutes prior to the TRV opening. The TRV was open for less than 1 minute, once power to the plant had been restored. The proper authorities were notified of the event.

During the time frame of this report, 10 drums were excavated, sampled, tagged, and over packed. The over packed drums are currently in temporary storage within the exclusion zone. The contractor is awaiting finger print lab analyses (characterization) of the drums in order to safely group or bulk the drums prior to shipment off-site for final disposition. Packaged drum carcasses and used PPE were shipped to a hazardous waste landfill.

Related site activities during this period included:

* Weekly progress meeting held on Tuesday, June 2, 1998. * Two options for the final site grading (full size plans and cross-sections) were given to the USEPA for discussion and decision by the City of Lock Haven and PADEP.

Major activities for the next week include: Operating the TDF, excavation, soil feed preparation, off-site and perimeter air monitoring, WWTP operations, ash handling, rock & inorganic debris washing, and drum handling.

PREPARED BY: Steve Gibson for Bill Werntges, USACE

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.